\begin{thebibliography}{99}

\bibitem{BNL} H.N. Brown et al., Muon $(g-2)$ Coll.,
Phys. Rev. Let. {\bf 86} (2001) 2227; B.L. Roberts,
these proceedings

\bibitem{CERN} J. Bailey et al., Nucl. Phys. B {\bf 150} (1979) 1

\bibitem{SCH48} J. Schwinger, Phys. Rev. {\bf 73} (1948) 416;
Phys. Rev. {\bf 76} (1948) 790

\bibitem{KM90} T. Kinoshita and W.J. Marciano,
in {\em Quantum Electrodynamics}, T. Kinoshita (ed), 
World Scientific, Singapore, (1990) p. 419

\bibitem{CM99} A. Czarnecki and W.J. Marciano,
 Nucl. Phys. B (Proc. Suppl.) {\bf 76} (1999) 245

\bibitem{MT00} P.J. Mohr and B.N. Taylor, 
Rev. Mod. Phys. {\bf 72} (2000) 351

%\bibitem{QED} C.M. Sommerfield, Phys. Rev. {\bf 107} (1957) 328;
%Ann. Phys. (N.Y.) {\bf 5} (1958) 26; A. Petermann, Helv. Phys. Acta
%{\bf 30} (1957) 407; Nucl. Phys. B {\bf 5} (1958)  677;
%H.H. Elend, Phys. Lett. B {\bf 20} (1966) 682; Erratum: ibid.
%{\bf 21} (1966) 720; B. Lautrup and E. de Rafael, Nucl. Phys. B
%{\bf 70} (1974) 317; Erratum: ibid. {\bf 78} (1974) 576;
%C. Chlouber and M.A. Samuel, Phys. Rev. D {\bf 16} (1977) 3596;
%A.S. Yelkhovsky, Sov. J. Nucl. Phys. {\bf 49} (1989) 656;
%A.I. Milstein and A.S. A.S. Yelkhovsky, 
%Phys. Lett. B {\bf 233} (1989) 11;
%T. Kinoshita, B. Ni\u{z}i\'c, and Y. Okamoto, 
%Phys. Rev. D {\bf 41} (1990) 593;
%R.Z. Roskies, E. Remiddi, and M.J. Levine, in {\em Quantum 
%Electrodynamics}, T. Kinoshita (ed), World Scientific, 
%Singapore, (1990) p. 162; M.A. Samuel and G. Li, Phys. Rev. D 
%{\bf 44} (1991) 3935;  Erratum: ibid. {\bf 48} (1991) 1879; 
%S. Laporta and E. Remiddi, Phys. Lett. B {\bf 265} (1991) 182; 
%ibid. {\bf 301} (1993) 440; ibid. {\bf 356} (1995) 390; 
%ibid. {\bf 379} (1996) 283; A.L. Kataev, Phys. Lett. B 
%{\bf 284} (1992) 401; T. Kinoshita, Phys. Rev. D 
%{\bf 47} (1993) 5013; Phys. Rev. Lett. {\bf 75} (1995) 4728; 
%Rep. Prog. Phys. {\bf 59} (1996) 1459; IEEE Trans. Instrum. Meas.
%{\bf 46} (1997) 108; S. Laporta, Phys. Lett. B {\bf 312} (1993) 495; 
%Nuovo Cim. A {\bf 106} (1993) 675; Phys. Rev. D {\bf 47}
%(1993) 4793; Phys. Lett. B {\bf 328} (1994) 522;
%ibid. {\bf 343} (1995) 421;  
%G. Li, R. Mendel, and M.A. Samuel, Phys. Rev. A {\bf 47} (1993) 1723;
%P.A. Baikov and D.J. Broadhurst, Proc. of {\em New Computing
%Techniques in Physics Research}, (1995), p. 167;
%S.G. Karshenboim, Phys. Atom. Nucl. {\bf 56} (1993) 857; 
%D.J. Broadhurst, Z. Phys. C {\bf 58} (1993) 339; 
%T. Kinoshita and M. Nio, Phys. Rev. Lett. {\bf 72} (1994) 3803; 
%Phys. Rev. D {\bf 53} (1996) 4909; 
%A.L. Kataev and V.V. Starshenko, Phys. Rev. D {\bf 52} (1995) 402;
%J. Ellis, M. Karliner, M.A. Samuel, and E. Seinfelds, hep-ph/9409376; 
%A. Czarnecki and M. Skrzypek, Phys. Lett. B {\bf 449} (1999) 354;
%V.W. Hughes and T. Kinoshita, Rev. Mod. Phys. {\bf 71} (1999) S133;  

\bibitem{oneloopEW} 
S.J. Brodsky and J.D. Sullivan, Phys. Rev. {\bf 156} (1967) 1644;
R. Jackiw and S. Weinberg, Phys. Rev. D {\bf 5} (1972) 2396;
I. Bars and M. Yoshimura, Phys. Rev. D {\bf 6} (1972) 374;
K. Fujikawa, B.W. Lee, and A.I. Sanda, Phys. Rev. D {\bf 6} 
(1972) 2923;
G. Altarelli, N. Cabibbo, and L. Maiani, Phys. Lett. B {\bf 40} 
(1972) 415;
W.A. Bardeen, R. Gastmans, and B.E. Lautrup, Nucl. Phys. B 
{\bf 46} (1972) 315 

\bibitem{twoloopEW} T.V. Kuhto, E.A. Kuraev, A. Schiller,
and Z.K. Silagadze, Nucl. Phys. B {\bf 371} (1992) 567;
S. Peris, M. Perrottet, and E. de Rafael,
Phys. Lett. B {\bf 355} (1995) 523;
A. Czarnecki, B. Krause, and W.J. Marciano,
Phys. Rev. D {\bf 52} (1995) R2619; 
Phys. Rev. Lett. {\bf 76} (1996) 3267

\bibitem{PDG00} Review of Particle Physics, D.E. Groom et al., 
Eur. Phys. J. C {\bf 15} (2000) 1

\bibitem{DG98} G. Degrassi and G.F. Giudice, Phys. Rev. D {\bf 58}
(1998) 053007

\bibitem{dRA94} E. de Rafael, Phys. Lett. B {\bf 322} (1994) 239

\bibitem{BPP96} J. Bijnens, E. Pallante, and J. Prades, 
Nucl. Phys. B {\bf 474} (1996) 379; Erratum: to be published; 
Phys. Rev. Lett. {\bf 75} (1995)
1447; Erratum: ibid. {\bf 75} (1995) 3781

\bibitem{HKS96} M. Hayakawa and T. Kinoshita,
Phys. Rev. D {\bf 57} (1998) 465; hep-ph/0112102;
M. Hayakawa, T. Kinoshita,  
and A.I. Sanda,  Phys. Rev. D {\bf 54} (1996) 3137; 
Phys. Rev. Lett. {\bf 75} (1995) 790

\bibitem{KN01} M. Knecht and A. Nyffeler, hep-ph/0111058

\bibitem{BAR01} E. Bartos et al., hep-ph/0106084

\bibitem{KNPR01} M. Knecht, A. Nyffeler, M. Perrottet, 
 and E. de Rafael, hep-ph/0111059

\bibitem{BCM01} I. Blokland, A. Czarnecki, and K. Melnikov, 
    hep-ph/0112117

\bibitem{ENJL} J. Bijnens, C. Bruno, and E. de Rafael, 
Nucl. Phys. B. {\bf 390} (1993) 501; 
J. Bijnens, E. de Rafael, and H. Zheng, Z. Phys. C {\bf 62} (1994) 437;
J. Bijnens and J. Prades, Phys. Lett. B {\bf 320} (1994) 130;
 Z. Phys. C {\bf 64} (1994) 475;  
J. Prades, Z. Phys. C {\bf 63} (1994) 491; 
J. Bijnens, Phys. Rep. {\bf 265} (1996) 369

\bibitem{PMR98} S. Peris, M. Perrottet, and E. de Rafael,
J. High Energy Phys. {\bf 05} (1998) 011

\bibitem{CLEO} J. Gronberg et al., CLEO Coll., 
Phys. Rev. D {\bf 57} (1998) 33; V. Savinov, hep-ex/9507005 

\bibitem{pi0form} 
H.-J. Behrend et al.,  CELLO Coll., Z. Phys. C {\bf 49} (1991) 401;
H. Aihara et al., TPC/2$\gamma$ Coll., Phys. Rev. Lett. {\bf 64} 
(1990) 172; 
C. Berger et al., PLUTO Coll., Phys. Lett. B {\bf 142} (1984) 125;
R.I. Djhelyadin et al., Lepton-G Exp., Phys. Lett. B {\bf 94} (1980) 548

\bibitem{BP01} J. Bijnens and F. Persson, hep-ph/0106130

\bibitem{BR68} S.J. Brodsky and E. de Rafael, 
Phys. Rev. {\bf 168} (1968) 1620; M. Gourdin and E. de Rafael, 
Nucl. Phys. B {\bf 10} (1969) 667

%\bibitem{oldvacpol} 
%C. Bouchiat and L. Michel, J. Phys. Radium {\bf 22}
%(1961) 121; L. Durand III, Phys. Rev. {\bf 128} (1962) 441;
%T. Kinoshita and R.J. Oakes, Phys. Lett. B {\bf 25} (1967) 143; 
%J.E. Bowcock, Z. Phys. {\bf 211} (1968) 400;
%M. Gourdin and E. de Rafael, Nucl. Phys. B {\bf 10} (1969) 667;
%A. Bram\`on, E. Etim, and M. Greco, Phys. Lett. B {\bf 39} (1972) 
%514; J. Bailey et al., Phys. Lett. B {\bf 55} (1975) 420;
%V. Barger, W.F. Long, and M.G. Olsson, Phys. Lett. B {\bf 60}
%(1975) 89; J. Calmet, S. Narison, M. Perrottet, and E.
%de Rafael, Phys. Lett. B {\bf 61} (1976) 283; 
%Rev. Mod. Phys. {\bf 49} (1977) 283; S. Narison, J. Phys. G {\bf 4}
%(1978) 1849; L.M. Kurdadze et al., Sov. J. Nucl. Phys.
%{\bf 40} (1984) 286; T. Kinoshita, B. Ni\u{z}i\'c, and Y. Okamoto, 
%Phys. Rev. Lett. {\bf 52} (1984) 717; Phys. Rev. D {\bf 31} (1985) 2108;
%L.M. Barkov et al., Nucl. Phys. B {\bf 256} (1985) 365;
%J.A. Casas, C. L\'opez, and F.J. Yndur\'ain, Phys. Rev. D {\bf 32}
%(1985) 736; \u{L}. Martinovi\u{c} and S. Dubni\u{c}ka,
%Phys. Rev. D {\bf 42} (1990) 884; A.Z. Dubni\u{c}kov\'a,
%S. Dubni\u{c}ka, and P. Strizenec, Acta Phys. Slov. {\bf 45} 
%(1995) 467; G. Cveti\u{c}, T. Lee, and I. Schmidt, hep-ph/0107069
 
\bibitem{vacpol}
S. Eidelman  and F. Jegerlehner,  Z. Phys. C {\bf 67} (1995) 585; 
K. Adel and F.J. Yndur\'ain, 
Rev. Acad. Ciencias (Esp) (1995), hep-ph/9509378;
D.H. Brown, W.A. Worstell, Phys. Rev. D {\bf 54} (1996) 3237; 
F. Jegerlehner, Nucl. Phys. B (Proc. Suppl.) {\bf 51C} (1996) 131;
 hep-ph/9901386 

\bibitem{MR01} W.J. Marciano and B.L. Roberts, hep-ph/0105056

\bibitem{eedata} R.R. Akhmetsin et al., Preprint Budker INP 99-10,
hep-ex/9904027; J.Z. Bai et al., BES Coll., hep-ex/0102003

\bibitem{JEG01} F. Jegerlehner, hep-ph/0104304

\bibitem{TY01} J.F. de Troc\'oniz and F.J. Yndur\'ain, hep-ph/0106025

\bibitem{KUE99}S. Binner, J.H. K\"uhn, and K. Melnikov, Phys. Lett. B
{\bf 459} (1999)  279; H. Czyz and J.H. K\"uhn, Eur. Phys. J. C
{\bf 10} (2001) 497; 
G. Rodrigo, A. Gehrmann-De Ridder, M.Guilleaume, and J.H. K\"uhn,
Eur. Phys. J. C {\bf 22} (2001) 81

\bibitem{HGJ01} A. H\"ofer, J. Gluza, and F. Jegerlehner, hep-ph/0107154

\bibitem{taudata}  R. Barate et al., ALEPH Coll., 
Z. Phys. C {\bf 76} (1997) 15; Eur. Phys. J. C {\bf 4} (1998) 409;
K. Ackerstaff et al., OPAL Coll.,  Eur. Phys. J. C {\bf 7} (1999) 571;
S. Anderson et al., CLEO Coll., Phys. Rev. D  {\bf 61} (2000) 112002

\bibitem{ADH98} R. Alemany, M. Davier, and A. H\"ocker, 
Eur. Phys. J. C {\bf 2} (1998) 123

\bibitem{DH01} A. H\"ocker, Talk at {\em XXXVI  Rencontres de Moriond,
ElectroWeak Interactions and Unified Theories},
 Les Arcs, France, March 10-17, 2001;
M. Davier, Nucl. Phys. B (Proc. Suppl.) {\bf 76} (1999) 327; 
hep-ex/9912044;
M. Davier and A. H\"ocker, Phys. Lett. B {\bf 419} (1998)
419; ibid. {\bf 435} (1998) 427

\bibitem{CEN01} V. Cirigliano, G. Ecker, and H. Neufeld, 
Phys. Lett. B {\bf 513} (2001) 361

\bibitem{NAR01} S. Narison, Phys. Lett. B {\bf 513} (2001) 53 

\bibitem{MEL01} K. Melnikov, Int. J. Mod. Phys. A {\bf 16} (2001) 4591

\bibitem{KRA97} B. Krause, Phys. Lett. B {\bf 390} (1997) 392

\bibitem{EID01} S.I. Eidelman, Nucl. Phys. B (Proc. Suppl.)
{\bf 98} (2001) 281

\bibitem{PP01} A. Pich and J. Portol\'es, Phys. Rev D {\bf 63} (2001)
093005

\end{thebibliography}

\documentstyle[12pt,a4,epsf]{article}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ba}{\begin{eqnarray}}
\newcommand{\ea}{\end{eqnarray}}

\begin{document}
\begin{titlepage}
\begin{flushright}
{Revised}\\
{CAFPE-1/01}\\
{UG-FT-132/01}\\
\end{flushright}
\vspace{2cm}
\begin{center}

{\large\bf 
The Standard Model Prediction for Muon $g-2$
 \footnote{Invited talk at ``Kaon 2001'', 6-13 June 2001, Pisa, Italy.
Work supported in part by
 the European Union TMR Network $EURODAPHNE$ (Contract
No. ERBFMX-CT98-0169), by MCYT, Spain (Grant No. FPA2000-1558),
and by Junta de Andaluc\'{\i}a, (Grant No. FQM-101).}}\\
\vfill
{\bf Joaquim Prades}\\[0.5cm]
Centro Andaluz de F\'{\i}sica de las Part\'{\i}culas
Elementales (CAFPE) and Departamento de
 F\'{\i}sica Te\'orica y del Cosmos, Universidad de Granada\\
Campus de Fuente Nueva, E-18002 Granada, Spain\\[0.5cm]
\end{center}
\vfill
\begin{abstract}
\noindent
The present Standard Model prediction for muon $g-2$
is reviewed. Emphasis is put in discussing the
main hadronic uncertainties.
\end{abstract}
\vfill
December 2001
\end{titlepage}

\section{Introduction}
Recently, the Muon $g-2$ Collaboration from 
the E821 experiment at Brookhaven National Lab (BNL)
 \cite{BNL} reported a new result 
for the  muon $g-2$ with an uncertainty more than five
times smaller than the last CERN
experiment \cite{CERN}. The E821 result 
when combined with the previous
experiments produce the present world average
\be
\label{BNL}
a_\mu\equiv \frac{|g_\mu|-2}{2} =
 ( 11 \ 659 \ 202.3 \pm 15.1 ) \cdot 10^{-10} \, .
\ee
The expected impressive final 
goal of E821 is to achieve an experimental
uncertainty in $a_\mu$ of the order of  $4 \cdot 10^{-10}$.

Accompanying this great experimental performance
a lot of effort has been put in the theoretical
side  to get a Standard Model prediction
for this quantity since the pioneering work 
of Schwinger \cite{SCH48}. In the next Sections I will review
the present Standard Model
prediction putting emphasis in discussing the 
main hadronic uncertainties which at present are the dominant.

\section{Quantum Electro-Weak-Dynamics Contributions}
\subsection{QED Contribution}
The QED contribution to muon $g-2$ is known to $O(\alpha^5)$
\ba
a_\mu^{\rm QED}&=& \frac{1}{2} \, \left(\frac{\alpha}{\pi}\right) +
0.765 \ 857 \ 376 (27) \, \left(\frac{\alpha}{\pi}\right)^2 +
24.050 \, 508 \ 98 (44) \, \left(\frac{\alpha}{\pi}\right)^3 
\nonumber \\ &+&
126.07 (41) \, \left(\frac{\alpha}{\pi}\right)^4 +
930 (170) \, \left(\frac{\alpha}{\pi}\right)^ 5 \, + \cdots 
\ea
Higher orders are negligible compared with the experimental 
uncertainty. The original references leading to this result 
can be found in \cite{KM90,CM99,MT00}.
The first, second, and third orders are known fully analytically.
The fourth order contains contributions 
which are only known numerically \footnote{Recently, 
T. Kinoshita has found a numerical error in the fourth order
 contribution which final result is not public yet.
However,  the total  Standard Model uncertainty is not upset by it.}. 
The fifth order is a numerical estimate of the dominant diagrams
enough for the BNL expected uncertainty.


Using the present world average value 
$\alpha^{-1}=137.035 \ 999 \ 76 (50)$ \cite{MT00},
one gets
\be
\label{aQED}
a_\mu^{\rm QED}=(11 \ 658 \ 470.6 \pm 0.3) \cdot 10^{-10} \, . 
\ee
which gives the bulk of the experimental value of $a_\mu$.

\subsection{Electroweak Contribution}
This contribution can be written as
\be
a_\mu^{\rm EW}= \frac{5}{3} \frac{G_F m_\mu^2}{8 \pi^2 \sqrt 2}
\left[ 1 + \frac{1}{5} 
\left( 1-4 \sin_{\theta_W}^2\right)^2 + (-97.0\pm 8.8)
\left(\frac{\alpha}{\pi} \right)+ \cdots \right] \, 
\ee
where the first term is the one-loop contribution \cite{oneloopEW}
and the second term is the $O(G_F m_\mu^2 \alpha)$ 
two-loop contribution \cite{twoloopEW}. 
Using \cite{PDG00}, one gets
\be
\label{aEW}
a^{\rm EW}_\mu= ( 19.4808\cdots +(- 4.4 \pm 0.4)) \cdot 10^{-10}
= (15.1 \pm 0.4)\cdot 10^{-10} \, 
\ee
for the one-loop plus two-loop EW contributions.
The uncertainty in the two-loop result 
takes into account uncertainties
in the Higgs mass, quark two-loop effects and 
$\alpha^3$ and higher corrections.
The leading $\ln (M_Z/m_\mu)$ logs can
be re-summed to all orders in $\alpha$ in  \cite{DG98}
and the result enlarge slightly the two-loop number
but it is well within the quoted uncertainty, so that
we keep the full two-loop result.

\section{Hadronic Contributions}
\subsection{Hadronic Light-by-Light Contribution}

This is the contribution  of
a hadronic Green's function with four legs coupled
to  electromagnetic two-quark currents.
This four-point function is attached in all possible ways 
with three of its legs to the muon line.
This contribution is of $O(\alpha^3)$ and 
cannot be related to any measured quantity
and we have to rely in our ability of treating
the strong interactions at {\em all} energies.

There are always {\em two} 
topologies to a full four-point function; namely,
first, two three-point form factors joined with full 
propagators, these we call three-point-like
contributions, and second, the pure four-point form factor
 which we refer to as the four-point-like contribution. 
In {\em all} cases the three vector legs are joined 
to the muon line through full vector two-point functions.  
We calculate the  leading $O(N_c)$ contributions to the full
four-point function as well as NLO in $1/N_c$  corrections 
which are saturated by four-point-like charged pion and kaon loops.
Being of different order in $1/N_c$, they {\em do not} have to match
the leading order in $1/N_c$ contributions as the
quark-loop contribution \cite{dRA94,BPP96}.

Based in the $1/N_c$ analysis of \cite{dRA94},
there have been two full calculations of this contribution
in \cite{BPP96} and \cite{HKS96}. 
Here, I pay more attention
to \cite{BPP96} and a full comparison with \cite{HKS96}
and references to previous work can be found in \cite{BPP96}.

Recently, there have also calculations of the pseudo-scalar
exchange in \cite{KN01} and of the pseudo-scalar
and scalar exchanges in \cite{BAR01}.
In \cite{KNPR01}, the coefficient of the leading  divergent logarithm 
has been obtained  analytically for the point-like Wess-Zumino-Witten
$\pi^0\gamma\gamma$ form factor. This result has been 
confirmed in \cite{BCM01}.

The results we get are\cite{BPP96}\footnote{Notice that we
 correct for an 
errata in the  sign of the pseudo-scalar and axial-vector
exchange contributions.}
\ba
\label{BPP}
a_\mu^{L-b-L}&=&((2.1\pm0.3)+(-0.68\pm0.2)
+(8.5\pm1.3)+(0.25\pm0.1))\cdot 10^{-10}\nonumber \\
&+&(-1.9\pm0.5)\cdot 10^{-10}=((3.8\pm0.3)+(4.4\pm2.1))
\cdot 10^{-10}\nonumber \\ &=& (8.3\pm3.2) \cdot 10^{-10} \,, 
\ea
where  the first term is the four-point-like contribution,
the second, third, and fourth terms are the three
types of three-point-like contributions, namely, 
when the propagators are scalar, pseudo-scalar, or axial-vector. 
The first term in the second line is the NLO in $1/N_c$ contribution.
The second split gives the contributions below 0.5 GeV and above 
0.5 GeV, respectively. 

With the same split as (\ref{BPP}),  
the result obtained in \cite{HKS96} is
\ba
\label{HKS}
a_\mu^{L-b-L}&=&((0.97\pm 1.1)+(0.0\pm0.0)
+(8.27\pm0.60)+(0.17\pm0.1))\cdot 10^{-10}\nonumber \\
&+&(-0.45\pm0.8))\cdot 10^{-10}
=(8.96\pm1.54) \cdot 10^{-10} \,.
\ea
 Two main features appear from both results.
The dominant contribution by far is the pseudo-scalar
exchange and second, there is a very  good
agreement  in  the contribution from the pseudo-scalar
exchange while for the rest of the $O(N_c)$ contributions
the disagreement is pretty  small. This disagreement
is larger in the NLO in $1/N_c$ contribution.
In fact both contributions
cancel out very much and the full result  is in quite
good agreement. Further scrutiny  of this cancellation 
in a  as much as possible model independent way is needed.

 For the low energy part (below 0.5 GeV)
of the four-point-like  and  scalar
exchange three-point-like contributions --they are related by Ward
identities-- we use the ENJL model \cite{ENJL}.
 For the higher energy part of the four-point-like
contribution we use a heavy quark
which mass acts as an IR cut-off \cite{BPP96}. The ENJL model has
{\em not} free parameters, they are fixed to low-energy $\pi\pi$ data.
In the ENJL model, 
full vector two-point functions have an nonphysical behavior
 \cite{BPP96,PMR98}
at intermediate energies which we corrected in this work.

The pseudo-scalar three-point-like contribution is
dominated by the $\pi^0$ exchange with non-negligible
contributions of the $\eta$ and $\eta'$ exchanges \cite{BPP96}.
We used a variety of $\pi^0\gamma^*\gamma^*$ form factors 
fulfilling all the known constraints. Fortunately, this
form factor is very constrained by the U$_A$(1) anomaly
which gives its normalization at the origin and by $\rho^0 \to
\pi^0 \gamma$ which gives the slope at the origin.
There are also  data on  $\pi^0 \to \gamma^* \gamma$ between 0.5 GeV and
3.3 GeV \cite{CLEO,pi0form}.
All these constraints make the model dependence 
small and it is also the reason of the good agreement of 
(\ref{BPP}) and (\ref{HKS}).
Nevertheless, the uncertainty in this contribution
can be reduced using data on $e^+e^-\to \pi^0 e^+ e^-$
at intermediate energies \cite{BP01}. Data
below 0.5 GeV on $\pi^0\to \gamma^* \gamma$ can also help.

The authors of \cite{KN01} were able to obtain
analytical formulas for the pseudo-scalar exchange
for a general class of $\pi^0 \gamma^* \gamma^*$ form factors
that fulfill the OPE and  large $N_c$ QCD constraints,
and which are  compatible with the data. Their result 
for this contribution $(8.3 \pm 1.2) \cdot 10^{-10}$ agrees
very well with both \cite{BPP96} and \cite{HKS96}
after correcting for the sign mistake.

For the NLO in $1/N_c$ contribution we cannot use the ENJL model.
We saturate them with charged pion and kaon loops coupled to photons
and need $P^+ P^- \gamma^* \gamma^*$ and $P^+ P^- \gamma^*$
vertices for which we take complete VMD  which works
very well for one-photon couplings at all energies. For two-photons
there is no data beyond $\pi^+ \pi^- \to \gamma \gamma$.  
In \cite{HKS96}, a HGS symmetry model was used and the discrepancy 
with \cite{BPP96} already
at low-energy is large but being of CHPT order $p^6$ 
there is no data to disentangle it. It is possible that information
on $e^+e^-\to\pi^+\pi^-e^+e^-$ could help to eliminate this
large model dependence.

The uncertainty in (\ref{BPP}) is obtained by adding {\em linearly}
the individual errors plus  $0.8 \cdot 10^{-10}$ added 
linearly to take into account the discrepancy between
our result and the one in (\ref{HKS}).
The uncertainty in (\ref{HKS}) is obtained adding quadratically their
individual errors.

I take the average of (\ref{BPP}) and (\ref{HKS}) 
\be
\label{aLbL}
a_\mu^{L-b-L}=(8.6\pm3.2) \cdot 10^{-10} \,
\ee
as the present value for this hadronic light-by-light
contribution. The uncertainty is the one from (\ref{BPP}) since already
takes into account the discrepancy between both results as explained
above and is more realistic. Improving both in the 
$\pi^0\gamma^*\gamma^*$ and $\pi^+ \pi^- \gamma^* \gamma^*$ vertices
could decrease this uncertainty to around $2\cdot 10^{-10}$.

\subsection{Hadronic Vacuum Polarization Contribution}
This contribution starts at order $\alpha^2$ and it is the
one with the largest uncertainty at present.
It is a hadronic Green's function with two legs coupled
to  electromagnetic two-quark currents and attached to the muon line. 

The $O(\alpha^2)$ contribution  can be written as \cite{BR68}, 
\be
\label{integral}
a_\mu^{(2){\rm hvp}}=\int^\infty_{4 m_\pi^2} {\rm d} t \ K(t) \, 
\frac{\sigma^{(0)}(e^+e^-\to {\rm hadrons})(t)}
{\sigma^{(0)}(e^+e^-\to\mu^+\mu^-)}
\ee
where $K(t)$ is a know function of $t$. There have been
many calculations with increasing accuracy due to better data
and theoretical input, see \cite{vacpol} for recent calculations
and \cite{MR01} for a critical review.
References to previous work can be found in \cite{KM90,CM99,MT00,MR01}.
 Since its uncertainty can be reduced 
systematically with accurate data one should consider this 
contribution  mainly of experimental origin.

The most recent $e^+e^-$ data \cite{eedata} from BES-II at Beijing,
and SND and CMD-2 at Novosibirsk has been used in \cite{JEG01}
\be
a_\mu^{e^+e^- \rm hvp}
=(697.4\pm10.5)\cdot 10^{-10} \, .
\ee
Also just using $e^+e^-$ data \cite{TY01} gets
\be
\label{TYee}
a_\mu^{e^+e^- \rm hvp}
=(698.2\pm9.7)\cdot 10^{-10} \, .
\ee 
These results give the  present average for this contribution  
\be
\label{eehad}
a_\mu^{e^+e^- \rm hvp} =(697.8\pm10.5)\cdot 10^{-10} \, .  
\ee
Further reduction of the uncertainty in the $e^+e^-$ data 
below the tau mass to the order or below 1\% 
is expected from VEPP-2M (CMD-2, SND) at Novosibirsk, 
DA$\Phi$NE (KLOE) at Frascati, BEPC (BES) at Beijing and other 
low energy  facilities. In the theory side, the possibility of measuring
the pion form factor at DA$\Phi$NE  and discussion of the relevant
QED corrections was presented in \cite{KUE99}.
The complete $O(\alpha)$ 
QED initial state, final state, and initial-final state radiation 
corrections to $e^+e^-\to \pi^+\pi^-$ has been recently presented
\cite{HGJ01}. 

Using CVC, one can relate the $\tau^- \to \pi^-\pi^0\nu_\tau$
with 
$I=1$ vector channel data to the $e^+e^- \to \pi^+\pi^-$ data.
The very precise tau hadronic decay data \cite{taudata} from 
ALEPH and OPAL at CERN and CLEO at CESR 
 when supplemented with the SU(2) breaking corrections can help
in increasing the accuracy of the  $e^+e^-\to \pi^+\pi^-$
data which gives 70\% of the contribution to $a_\mu^{\rm hvp}$
and 80\% of its uncertainty
when integrated in (\ref{integral}) from $4m_\pi^2$ to $t=0.8$ GeV.
This program was started in \cite{ADH98} 
and continued in \cite{DH01} where pQCD was also
pushed down up to 1.8 GeV helping to reduce the final uncertainty 
sizeably. This use has been recently confirmed
by the BES-II data which give compatible results \cite{TY01,DH01}.

Adding the rest of  the hadronic vacuum 
polarization contributions and the known isospin breaking corrections, 
\cite{DH01} gets
\be
\label{DHtau}
a_\mu^{\tau \,  \rm hvp}
=(692.4\pm6.2)\cdot 10^{-10} \, .
\ee
The leading isospin breaking corrections studied in \cite{CEN01} 
agree quite well within errors with the same corrections 
applied in \cite{DH01}.
Numerically, the isospin corrections applied in \cite{TY01}
to the tau data are very similar to the ones in \cite{DH01}.
The result from tau data in \cite{TY01} is
\be 
\label{TYtau}
a_\mu^{\tau \, \rm hvp}
=(695.2\pm6.4)\cdot 10^{-10} \, .
\ee 
From these results the present average 
for the $a_\mu^{\rm hvp}$ from tau data reads
\be
\label{tauhad}
a_\mu^{\tau \, \rm  hvp} =(693.8\pm6.4)\cdot 10^{-10} \, . 
\ee
The recent calculation \cite{NAR01} is also compatible within errors
though no isospin breaking effects are included.
For the combined final result of (\ref{eehad}) and (\ref{tauhad}) 
I take the weighted average
\be
\label{had}
a_\mu^{(\alpha^2) \rm hvp} =(694.9\pm 6.4)\cdot 10^{-10} \, . 
\ee

There are some $O(\alpha^3)$ contributions already included in 
(\ref{had}), namely, the intermediate $\pi^0 \gamma^*$ and 
$\eta \gamma^*$ states in $\sigma(e^+e^-\to {\rm hadrons})$.
The rest of the $O(\alpha^3)$   corrections 
have been calculated in \cite{ADH98,KRA97} and are well
under control
\be
\label{alpha3}
a_\mu^{(\alpha^3) \, \rm hvp}= (-10.0\pm0.6)\cdot 10^{-10} \, .
\ee

It was pointed out in \cite{MEL01} that final state radiative
corrections are eliminated up to 80 \% in the Novosibirsk
data analysis of $\sigma(e^+e^- \to \gamma^* \to {\rm hadrons})$. 
They were estimated in \cite{TY01}
and added back to their final number (\ref{TYee}).
These $O(\alpha^3)$ corrections depend obviously on the
experimental set-up and on the particular analysis of the data.
In fact, they were already taken into account in the ALEPH tau 
data\footnote{I thank Michel Davier for informing me on this point.}
used by \cite{ADH98} and \cite{TY01} and therefore should {\em not}
be added back to the results (\ref{DHtau}) and (\ref{TYtau}).

Summing (\ref{had}) and (\ref{alpha3}), one gets  for
the total hadronic vacuum polarization contribution
\be
\label{ahvp}
a_\mu^{\rm hvp }= (684.9\pm6.4)\cdot 10^{-10}\, .
\ee
The expected $e^+e^-$ data accuracy
below 1 \% supplemented with  theoretical efforts like 
\cite{KUE99,HGJ01}
will reduce the uncertainty in  $a_\mu^{\rm hvp}$ from 
$e^+e^-$ up to the order of $6\cdot 10^{-10}$.
Joint works of tau and $e^+e^-$ groups \cite{DH01,EID01} 
are also announced and will reduce this uncertainty further.
Isospin breaking studies like \cite{CEN01} will 
help to take these corrections under better control.
Chiral symmetry can also help to reinforce
the accuracy of the $\pi\pi$ dominant contribution \cite{PP01}.

\section{Results and Summary}
Summing all the Standard Model contributions, (\ref{aQED}),
(\ref{aEW}), (\ref{aLbL}), and (\ref{ahvp}) to muon $g-2$, one
gets
\be
\label{SM}
a_\mu^{\rm SM}=  ( 11\, 659\, 179.2\pm 9.4) \cdot 10^{-10} \, .
\ee
where uncertainties of the QED, EW, and hadronic light-by-light
contributions  have been added linearly and afterwards added
quadratically to  the hadronic vacuum polarization uncertainty.

As a final result, we get
\be
a_\mu-a_\mu^{\rm SM}=(23.1\pm 16.9)\cdot 10^{-10}
\ee
i.e. there is at present  a bit more than  one
sigma of discrepancy.
The significance of this discrepancy could be largely enhanced  
by the aimed experimental uncertainty BNL goal of $4\cdot 10^{-10}$.
The announced improvements on 
the hadronic contributions are also very interesting
and can reduce the uncertainty of the muon $g-2$ Standard Model 
prediction to the order of $6\cdot 10^{-10}$ to $7\cdot 10^{-10}$.
The near future of muon $g-2$ reveals thus very exciting.

\section*{Acknowledgements}
It is a pleasure to  thank Hans Bijnens, Michel Davier, Andreas H\"ocker,
and Lee Roberts for comments and reading the manuscript.

\end{document}